The basic barometer can be used to measure the height of the building. If the barometer reading at the top & batton of the building is 730mm Hg & 755 mm Hg. Determine the height of the building. the density of air 1.18 kg per cubic meter & 13600 kg per cubic meter


1
Expert's answer
2021-05-05T07:45:42-0400

"\triangle P=P_{down}-P_{top}=(h_{down}-h_{top}) \rho _{Hg} \times g"

"\triangle P=(755-730) \times 10^{-3} \times 13600 \times 9.81"

"\triangle P=\triangle y \rho_{air}g"

"\triangle y = \frac{\triangle P}{\rho_{air}g}= \frac{(755-730) \times 10^{-3} \times 13600 \times 9.81}{1.18 \times 9.81}= 288.136 m"